OpenStudy (anonymous):

What is the equation of the line perpendicular to the line y=(4/9)x-2 that passes through the point (4,3)?

OpenStudy (texaschic101):

the slope in the given equation, y = 4/9x - 2, is 4/9. However, we need the perpendicular slope, also called the negative reciprocal slope. All that means is " flip " the slope given and change the sign. So the negative reciprocal of 4/9 is -9/4. See how I flipped the slope and changed the sign. Now we use y = mx + b slope(m) = - 9/4 (4,3) x = 4 and y = 3 now we sub y = mx + b 3 = -9/4(4) + b 3 = - 9 + b 3 + 9 = b 12 = b Therefore, your perpendicular equation is : y = -9/4x + 12 any questions ?

OpenStudy (anonymous):

I've got it now, thanks for the detailed explanation <3

OpenStudy (texaschic101):

no problem :)